OpenStudy (anonymous):

A seedman has seeds worth 50¢ per pound mixed with seeds worth 80¢ per pound. If he wishes to make a mixture of 30 pounds to sell at 75¢ per pound, how many pounds of 80 cent seed should he use?

OpenStudy (anonymous):

Write out steps used

OpenStudy (anonymous):

do you want me to answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well wait iam thinking

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well we have infinity of solutions

OpenStudy (anonymous):

look, if we said that the number of pound of 50 cent seed is x and that of 80 seeds is y then 50x+80y=30*75 so for every given value of x we have value of y for example if x=1 then the number of pound of 80 cent seeds is: y=220/8

OpenStudy (anonymous):

ok. thank you ! but thats not an answer lol

OpenStudy (anonymous):

why? it is not?

OpenStudy (anonymous):

a) 5 b) 20 c) 25

OpenStudy (anonymous):

well tell me what are these and from which country you are?

OpenStudy (anonymous):

25 pounds 75(30)=80x+50(30-x) 2250=80x+1500-50x 750=30x x=25

OpenStudy (anonymous):

Thaaaannnkkkkk yyyyooouuuuuu !!!!!!! shinigami1m